Determine the capacitance of a bakelite dielectric capacitor whose plate area

is 1.25x10-4 m2. The plates are separated by 5.0x10-4 m.


1
Expert's answer
2022-05-25T10:27:29-0400
"C=\frac{\epsilon_0 A}{d}"

"C=\frac{8.85*10^{-12}*1.25*10^{-4}}{5.0*10^{-4}}=2.2*10^{-12}\:\rm F"